Nitrogen 13 has 7 protons and a half life of 600 seconds. What is the activity in Ci after 88.367 minutes? The initial mass of the nitrogen 13 is 93.633 micrograms. The mass of N13 is 13.005799 g/mole.
Answer #1

From the given question,

half life(t1/2)=600 sec

decay constant(\lambda)=ln (2)/t1/2

=ln (2)/600

=1.15x 10-3

N0=(93.633*10-6)*6.022*1023/13.005799

=4.33 x 1018 atoms

N=N0e-\lambdat

N=4.33 x 1018 e-(1.15x 10^-3x88.367x60)

=9.74x 1015

A=\lambdaN

=1.15x 10-3x 9.74x 1015

=1.12 x 1013 Bq

=(1.12 x 1013)/(3.7×1010)

=302.7 Ci

Activity is 302.7 Ci